A potters wheel having a radius 0.45 m in a moment of inertia of 11.1 kg . m2 is rotating freely at 50 rev/min. The potter can stop the wheel in 5.0 s by pressung a wet rag against the r ams extertimg a radially inward forcw of 67 N find the effective cowfficent of kinetic friction between the wheel ans the wet rag.
Answer #1

ωo = 50 rev/min = 50 * 2π/60 = 5.24 rad/s

ωt = ωo + α*t

0 =5.24 + α * 5---> α = -5.24/5 = 1.05 rad/s^2

Torque:

τ = I * α

τ = r * f ---> f = friction = μ * F

--> r * μ * F = I * α

     0.45 * μ * 67 = 11.1 * 1.05

--> μ = 0.386

Kindly upvote :)